How do you know if a value is a solution for an inequality and how is this different from determing a valua is a solution to equation?
Found 2 solutions by oscargut, stanbon:
Answer by oscargut(686)

to check a solution in inequalities and equations you have to substitute the values
examples:
x=1 is a solution for 2x+3=5 (2(1)+3=5)
x=2 is not a solution for 3x-3=0 (3(2)-3=3)
x=3 is a solution for 3x-1<10 (3(3)-1=8<10)
x=4 is not a solution for 5x-1>32 (5(4)-1=19<32)

Answer by stanbon(48546)

The procedure is the same.
If you find a variable value or two
values for an inequality statement,
substitute those values for the variable(s)
and see if you get a true statement.
---------------------------------------
Example:
3x + 4y >6
y > (-3/4)x + (3/2)
----------------------
Question:
Is (0,0) part of the solution?
Substitute to get 0 > (-3/4)*0+(3/2)
0 > 3/2
That is false so (0,0) is not part of the solution.
-----------------
Is (0,5) part of the solution?
Substitute to get 5 > (-3/2)*0+(3/2)
5 > (3/2)
That is true so (0,5) is part of the solution.
========================
Cheers,
Stan H.